# Absence of bulk charge density wave order in the normal state of UTe2

### v2

@@ II.3 Survey for putative CDW order in bulk
We start with the search for resolution-limited charge-order peaks. We surveyed reciprocal space in search of signal at $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$, corresponding to CDW wave-vector of type $\bm{q}_{\mathrm{1}}$ in STM, shown by the grey sphere in Fig. 3(a). The letter $a$ denotes the specific momentum transfer, where the CDW peak was studied in our experiments. Such charge order would be detected when correlation lengths are larger than $2/f_{\mathrm{res}}$, where $f_{\mathrm{res}}$ denotes the width in momentum space of resolution-limited structural Bragg peaks. We therefore chose an incident energy of 3.37 keV, where resolution volumes are maximal.
To account for possible uncertainty in the CDW position, we perform a 3-dimensional grid search at positions near $\bm{q}_{\mathrm{1}}$. The mesh grid coordinates (colored spheres in Fig. 3(a)) are spaced 0.0018 r.l.u apart in $k$ and $l$, which is smaller than the instrument resolution (green ellipsoid in Fig. 3(a)) to maximize the chance of detecting a resolution-limited CDW peak. The mesh consists of $H$-scans at given $K_{0}$, $L_{0}$-coordinates indicated in Fig. 3. The step width, $\Delta H=3.6\cdot 10^{-4}\,\mathrm{\AA}^{-1}$, is smaller by a factor of two than the experimental resolution determined for a nearby specular Bragg peak (c.f. Supplementary Material Text 3).
Figure 3(c) shows the resulting scans, where the color of the plot corresponds to the $K_{0}$, $L_{0}$-coordinates in (b). The scans can be modeled by only a constant background of $0.23(1)$ cts/s with with averaged standard error from Poisson counting statistics of $\sigma=0.19$ cts/s. As a conservative upper limit of the maximum intensity due to charge modulations, we assume that one would be able to detect signatures that are three error bars above the background (in the following denoted “$3\sigma$”).
@@
Figure 4: Absence of diffuse charge-scattering peaks in resonant X-ray diffraction data. The three panels show momentum-space cuts through positions, where according to STM reports diffuse charge scattering peaks of a multi-component CDW above the superconducting transition were reported. Shown in (a) is an $H$ scan through $\bm{q}_{\mathrm{1}}$ ($H,2.5,2.5$) using 3.73 keV incident energy, in (b) through $\bm{q}_{\mathrm{1}}$ ($H,3.5,3.5$) using 4.95 keV incident energy, and in (c) through $\bm{q}_{\mathrm{2}}$ ($H,2.338,2.661$) using incident energy 4.95 keV. As explained in the text and derived from STM, a diffuse CDW signal may be expected to have full width have maximums in the interval $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{max}}$. The red circles present X-ray counts, where the error bars correspond to uncertainties arising from Poisson counting statistics. Points at zero intensity are here shown with vanishing error bars. The background was fitted by linear functions (blue thick line). The green line corresponds to a Gaussian peak with an amplitude of $3\sigma$ above the background, where $\sigma$ denotes the mean error bar of the X-ray counts.
Turning now to the second scenario, where CDW peaks have $\sim$ nm-sized short correlation lengths, we expand our search to look for signal at the corresponding bulk positions of $\bm{q}_{\mathrm{1}}$ and $\bm{q}_{\mathrm{2}}$. For wavevector $\bm{q}_{\mathrm{1}}$, data was collected in two Brillouin zones using two incident energies: $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$ at 3.73 keV and $\bm{Q}^{\mathrm{CDW}}_{1,b}=(-0.54,3.5,3.5)$ at 4.95 keV. Here $b$ labels another momentum-transfer, where we studied the respective wave-vector. For wavevector $\bm{q}_{\mathrm{2}}$, scans were taken in one Brillouin zone at $\bm{Q}^{\mathrm{CDW}}_{2,b}=(-0.430,2.338,2.661)$ with incident energy of 4.95 keV.
Our scan results at both incident energies and wavevectors are presented in Fig. 4. Panels (a) and (b) show scans at $\bm{q}_{\mathrm{1}}$ at 3.73 keV and 4.95 keV, respectively, and panel (c) shows scans at $\bm{q}_{\mathrm{2}}$ at 4.95 keV.
The background intensities were each fitted using a line with a constant slope and offset, shown in blue. The fit results are displayed above each panel. Mean error bars, $\sigma$, as obtained from Poisson counting statistics are given by (a) $0.17\,\mathrm{cts}/\mathrm{s}$, (b) $8.9\,\mathrm{cts}/\mathrm{s}$, and (c) $8.38\,\mathrm{cts}/\mathrm{s}$.
@@
Comparison of intensities lets us conclude that the maximum intensities of CDW peaks having short correlation lengths in bulk are at least eight orders of magnitude smaller than for the structural Bragg peaks.
II.4 Upper boundary on atomic displacements
Charge modulations probed by Thomson scattering of X-rays can arise either from spatial modulation of valence electrons or from periodic atomic displacements [26]. These two effects typically appear together. With periodic lattice displacements leading to a much stronger response in X-ray scattering, we consider them as indirect evidence of valence modulations [27]. Here, we will calculate the limits of atomic displacements in UTe2 based on the scattering intensities reported above.
Using standard expressions, displacements of atoms along a direction $\bm{d}$ away from the equilibrium position $\bm{r}_{i}$ may be written as $\bm{r}^{\prime}_{i}=\bm{r}_{i}+u\bm{\hat{d}}\cdot\sin(\bm{q}_{0}\cdot\bm{r}_{i})$. In diffraction, this leads to charge-order Bragg peaks around structural Bragg peaks $\bm{q}$, appearing in leading order at $\bm{G}\pm\bm{q}_{0}$.
The structure factor of the modulated crystal can accordingly be written as
$S(\bm{Q})=\sum_{\textbf{G}}\delta(\bm{Q}-\bm{G})+i\biggl{(}\frac{\bm{Q}\cdot\bm{u}}{2}\biggr{)}\sum_{\textbf{G}}\biggl{[}\delta(\bm{q}+\bm{q}_{0}-\bm{G})+\delta(\bm{q}-\bm{q}_{0}-\bm{G})\biggr{]}$ (1)
and is related to the measured quantities given that the structure factor is proportional to the integrated intensities of Bragg peaks, i.e., $I(\bm{Q})\propto|S(\bm{Q})|^{2}$. The ratio of intensities of a structural Bragg peak to a charge order Bragg peak yields
@@ II.4 Upper boundary on atomic displacements
The intensity at CDW position in momentum space may arise from a combination of two domains. For example, at momentum $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.545,2.5,2.5)$, the intensity arises from CDW modulations with wave-vector component $(-0.545,0.5,0.5)$ of the $\bm{G}_{1}=(0,2,2)$ zone and with $(-0.545,-0.5,-0.5)$ of the $\bm{G}_{2}=(0,3,3)$ zone, two different representatives of the $\bm{q}_{1}$-modulation. Calculation of the crystal structure factor of UTe2 permits comparison with the intensity of the structural Bragg peak at $\bm{G}_{0}=(0,1,1)$ and yields
$\left(\frac{\bm{u}\cdot\bm{Q}^{\mathrm{CDW}}_{0}}{2}\right)^{2}=\frac{A(\bm{Q}^{\mathrm{CDW}}_{1,a})\cdot V(\bm{Q}^{\mathrm{CDW}}_{1,a})}{A(\bm{G}_{0})\cdot V(\bm{G}_{0})}\cdot\frac{S(\bm{G}_{0})}{S(\bm{G}_{1})+S(\bm{G}_{2})}.$ (3)
@@
We first look at the measurements taken at 3.37 keV shown in Sec. II.3. For resolution-limited CDW peaks the maximum atomic displacements can be estimated using the above equation. Along the (100) direction we determine $u_{||(100)}\leq 4.7\cdot 10^{-3}\,\mathrm{\AA}$ and along the (110) direction $u_{||(011)}\leq 1.2\cdot 10^{-3}\,\mathrm{\AA}$. In comparison, the atomic displacements in the CDW in YBa2Cu3O6.67 is of order $u\approx 3\cdot 10^{-3}\,\mathrm{\AA}$ [27], or similarly $u\approx 10^{-3}\,\mathrm{\AA}$ in related compounds [28].
@@
Now we turn to the data shown in Fig. 4, where we have assumed short correlation lengths for the CDW. Again using Eq. 3 with consideration of multiple domains, we can find upper bounds for displacement amplitudes of CDWs having widths $f_{\mathrm{CDW}}=(f_{\mathrm{min}}+f_{\mathrm{max}})/2=0.09\,\mathrm{\AA}^{-1}$. Since the CDW peak width is much larger than the REXS momentum resolution, the integration volume is essentially resolution independent: $V(\bm{Q}_{\mathrm{CDW}})=(f_{\mathrm{CDW}}^{2}+f_{\mathrm{res}}^{2})^{3/2}\approx f_{\mathrm{CDW}}^{3}$.
Considering Fig. 4(b), we find that atomic displacements at 4.95 keV incident energy are limited by $u\leq 0.22\,\mathrm{\AA}$ along the $(100)$ direction and $u\leq 0.06\,\mathrm{\AA}$ along the $(011)$ direction. Considering instead Fig. 4(a), we find that for $\bm{q}_{\mathrm{I}}$ at incident energy of 3.73 keV, the atomic displacements are limited by $u<0.33\,\mathrm{\AA}$ along $(011)$ and $u\leq 1.2\,\mathrm{\AA}$ along the $(100)$ direction, therefore providing a weaker bound. While these numbers are larger than seen in the cuprates, the order in UTe2 for this case is localized and would represent inhomogeneity in the CDW state throughout the sample.
III Discussion
@@
Our measurements using resonant X-rays show that the charge modulations earlier observed in STM above the superconducting transition at the surface of UTe2 are absent from the bulk within our detection limits posed by state-of-the art X-ray diffraction. Our measurements were performed at $T=2.2$ K, about four times lower than the highest temperature the CDW in UTe2 is reported to exist [18]. Our results did not reveal diffraction intensity in the vicinity of the CDW wave vectors at either the U M4 or Te L1 absorption edges that is noticeable above the background level.
@@
The coverage of our fine mesh-grid search near the $\bm{q}_{1}$ CDW wave vector rules out the possibility that the CDW peak is too sharp to be observed in bulk X-ray measurements. Notably, the maximum diffraction intensity of long-ranged CDW is at least five orders of magnitude smaller relative to the intensity of the strongest Bragg peaks observed. Atomic displacements of long-range modulations with resolution limited Bragg signatures are four orders of magnitude smaller than the respective atomic units along $(100)$ and $(011)$.
@@
We investigated regions in reciprocal space large enough using step-widths fine enough in order to rule out the presence of CDW peaks with short correlation lengths in the bulk. Comparing REXS intensities, we find that integrated intensities of putative bulk CDW peaks are at least eight orders of magnitude weaker than nearby structural Bragg peaks. We conclude that in the case of putative broad CDW peaks, maximum atomic displacements are two orders of magnitude smaller than the respective crystallographic unit for both $(100)$ and $(011)$ reciprocal directions.
@@
We highlight that neither thermodynamic measurements [22, 23, 24] nor inelastic neutron scattering identified any signatures suggestive of bulk charge order above the superconducting transition [29, 30, 31, 32]. In combination with our results, the emergence of a surface charge-density wave in the normal state provides the most consistent picture. As for the superconducting state, our results point to two possible conclusions: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or these modulations are restricted to the surface. Bulk scattering measurements in the superconducting state are required to address this outstanding question.
IV Materials and Methods
@@
Resonant Elastic X-ray Scattering Experiments were carried out in the second experimental hutch EH2 of beamline P09 at the synchrotron source PETRA III [33]. The X-ray diffraction was carried out in a horizontal scattering geometry. The UTe2 sample was cooled using a variable temperature insert in a cryomagnet able to provide magnetic fields up to 14 T. X-rays with incident linear polarization were chosen using phase plates [34].
@@
where $A$ denotes the amplitude (or the maximum) and $E$ is a Gaussian profile with full width at half maxima $f_{\bm{d}_{1}}$, $f_{\bm{d}_{2}}$, and $f_{\bm{d}_{3}}$ along the three directions $\bm{d}_{1}=(1,0,0)$, $\bm{d}_{2}=(0,1,1)$, and $\bm{d}_{3}=(0,0,0)$, respectively. Integrated intensity of a Bragg peak is therefore given by $I=\frac{1}{8}\sqrt{\frac{\pi}{\ln(2)}}^{3}\cdot Af_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$. The profile of structural Bragg peaks is essentially restricted by resolution. We therefore define $f_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$ as the experimental resolution volume in momentum space.
Data Availability The datasets generated during and/or analysed during the current study are available from the corresponding author on request.

### v3

@@ VI.3 Survey for putative CDW order in bulk
We start with the search for resolution-limited charge-order peaks. Such long-range order has correlation lengths much larger than $2/f_{\mathrm{res}}$, where $f_{\mathrm{res}}$ denotes the width in momentum space of resolution-limited structural Bragg peaks. To increase our chance to discover charge-order and to avoid missing its signatures due to the high resolution of REXS, we chose an incident energy of 3.37 keV, where the resolution volume is larger compared to 4.95 keV.
We surveyed reciprocal space in the search of signal at $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$, corresponding to CDW wave-vector of type $\bm{q}_{1}$, shown by the grey sphere in the center of Fig. 3(A). The letter $a$ denotes the specific momentum transfer, where the CDW peak was studied in our experiments. To account for possible uncertainties in momentum transfers due to the accuracy of the diffractometer (see Supplementary Information Text S3), we performed a 3-dimensional grid search at positions near $\bm{Q}^{\mathrm{CDW}}_{1,a}$. The mesh consists of $H$-scans at given $K_{0}$, $L_{0}$-coordinates indicated in Fig. 3 in terms of spheres. The spacing along $K$ and $L$ was chosen as 0.0018 r.l.u apart, which is smaller than the instrument resolution (green ellipsoid in Fig. 3(A)). The step width in each linescan along $H$, $\Delta H=3.6\cdot 10^{-4}\,\mathrm{\AA}^{-1}$, is smaller by a factor of two than the experimental resolution determined for a nearby specular Bragg peak (c.f. Supplementary Material Text 3), and therefore small enough to account for uncertainties in the incommensurate component of $\bm{q}_{1}$ along $(100)$.
@@
Figure 4: Absence of diffuse charge-scattering peaks in resonant X-ray diffraction data. The three panels show momentum-space cuts through positions, where according to STM reports diffuse charge scattering peaks of a multi-component CDW above the superconducting transition were reported. Shown in (A) is an $H$ scan through $\bm{q}_{\mathrm{1}}$ at ($H,2.5,2.5$) using 3.73 keV incident energy, in (B) through $\bm{q}_{\mathrm{1}}$ at ($H,3.5,3.5$) using 4.95 keV incident energy, and in ((C)) through $\bm{q}_{\mathrm{2}}$ at ($H,2.338,2.661$) using incident energy 4.95 keV. As explained in the text and derived from STM, a diffuse CDW signal may be expected to have full width have maximums in the interval $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{max}}$. The red circles present X-ray counts, where the error bars correspond to uncertainties arising from Poisson counting statistics. Points at zero intensity are here shown with vanishing error bars. The background was fitted by linear functions (blue thick line). The green line corresponds to a Gaussian peak with an amplitude of $3\sigma$ above the background, where $\sigma$ denotes the mean error bar of the X-ray counts.
Turning now to the second scenario, where CDW peaks have $\sim$ nm-sized short correlation lengths, we expand our search to look for signal at the corresponding bulk positions of $\bm{q}_{\mathrm{1}}$ and $\bm{q}_{\mathrm{2}}$. For wavevector $\bm{q}_{\mathrm{1}}$, data were collected in two Brillouin zones using two incident energies: $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$ at 3.73 keV and $\bm{Q}^{\mathrm{CDW}}_{1,b}=(-0.54,3.5,3.5)$ at 4.95 keV. Here $b$ labels another momentum-transfer, where we studied the respective wave-vector. For wavevector $\bm{q}_{\mathrm{2}}$, scans were taken in one Brillouin zone at $\bm{Q}^{\mathrm{CDW}}_{2,b}=(-0.430,2.338,2.661)$ with incident energy of 4.95 keV.
The results of these scans for both incident energies and wavevectors are presented in Fig. 4. Panels (A) and (B) show scans at $\bm{q}_{\mathrm{1}}$ at 3.73 keV and 4.95 keV, respectively, and panel (C) shows scans at $\bm{q}_{\mathrm{2}}$ at 4.95 keV. The background intensities were each fitted using a line with a constant slope and offset, shown in blue. The fit results are displayed above each panel. Mean error bars, $\sigma$, as obtained from Poisson counting statistics are given by (A) $0.17\,\mathrm{cts}/\mathrm{s}$, (B) $8.9\,\mathrm{cts}/\mathrm{s}$, and (C) $8.38\,\mathrm{cts}/\mathrm{s}$.
@@
Comparison of intensities lets us conclude that the maximum intensities of CDW peaks having short correlation lengths in bulk on the Te L-edge are at least eight orders of magnitude smaller than for the structural Bragg peaks. Compared to that, in the cuprate YBa2Cu3O6.67 [37] even for non-resonant x-rays charge order signals are only seven orders of magnitude smaller relative to strong nearby structural Bragg peaks.
VI.4 Upper boundary on atomic displacements
Charge modulations probed by Thomson scattering of X-rays can arise either from spatial modulation of valence electrons or from periodic atomic displacements [38]. These two effects typically appear together and cause each other. With periodic lattice displacements leading to a much stronger response in X-ray scattering, we search for them directly and are further interested in them as indirect evidence of valence modulations [37]. Here, we will calculate the limits of atomic displacements in UTe2 based on the scattering intensities reported above.
Using standard expressions, harmonic displacements of atoms along a direction $\bm{d}$ away from the equilibrium position $\bm{R}_{i}$ may be written as $\bm{R}^{\prime}_{i}=\bm{R}_{i}+u\bm{\hat{d}}\cdot\sin(\bm{q}_{0}\cdot\bm{R}_{i})$. In diffraction, this leads to charge-order Bragg peaks around structural Bragg peaks $\bm{G}$, appearing in leading order at $\bm{G}\pm\bm{q}_{0}$.
Neglecting higher order terms, the structure factor of the modulated crystal around $\bm{Q}=\bm{G}$ can be written as
$S_{\mathrm{m}}(\bm{Q})=S(\bm{G})\delta(\bm{Q}-\bm{G})+S(\bm{G})\biggl{(}\frac{\bm{Q}\cdot\bm{u}}{2}\biggr{)}^{2}\delta(\bm{Q}-\bm{G}\pm\bm{q}_{0})\,.$ (1)
As the structure factor is proportional to the integrated intensities of Bragg peaks, i.e., $I(\bm{Q})\propto S(\bm{Q})$, the ratio of intensities of a structural Bragg peak to a charge order Bragg peak therefore yields for the amplitude $\bm{u}=u\bm{d}$ of the modulation around $\bm{G}$
@@ VI.4 Upper boundary on atomic displacements
The intensity at a given position in momentum space may arise from a combination of different charge-order satellites associated with neighboring Brillouin zones. For example, at momentum $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.545,2.5,2.5)$, the intensity arises from CDW modulations with wave-vector component $(-0.545,0.5,0.5)$ of the $\bm{G}_{1}=(0,2,2)$ zone and with $(-0.545,-0.5,-0.5)$ of the $\bm{G}_{2}=(0,3,3)$ zone, two different representatives of the $\bm{q}_{1}$-modulation. Calculation of the crystal structure factor of UTe2 (see Supplementary Material Text S5) permits comparison with the intensity of the structural Bragg peak at $\bm{G}_{0}=(0,1,1)$ and yields
$\left(\frac{\bm{u}\cdot\bm{Q}^{\mathrm{CDW}}_{1,a}}{2}\right)^{2}=\frac{A(\bm{Q}^{\mathrm{CDW}}_{1,a})\cdot V(\bm{Q}^{\mathrm{CDW}}_{1,a})}{A(\bm{G}_{0})\cdot V(\bm{G}_{0})}\cdot\frac{S(\bm{G}_{0})}{S(\bm{G}_{1})+S(\bm{G}_{2})}.$ (3)
@@
Following this procedure, we first look at the measurements taken at 3.37 keV shown in Sec. VI.3. For resolution-limited CDW peaks maximum atomic displacements can be estimated using the above equation for $V(\bm{Q}^{\mathrm{CDW}}_{1,a})\approx V(\bm{G}_{0})$. Along the (100) direction we determine $u_{||(100)}\leq 4.7\cdot 10^{-3}\,\mathrm{\AA}$ and along the (110) direction $u_{||(011)}\leq 1.2\cdot 10^{-3}\,\mathrm{\AA}$. In comparison, the atomic displacements in the CDW in YBa2Cu3O6.67 are of order $u\approx 3\cdot 10^{-3}\,\mathrm{\AA}$ [37], or similarly $u\approx 10^{-3}\,\mathrm{\AA}$ in related compounds [39].
@@
Now we turn to the data shown in Fig. 4, where we searched for long-range CDW order with nanometer-sized correlation lengths. Again using Eq. 3 with consideration of multiple domains, we can find upper bounds for displacement amplitudes of CDWs having widths $f_{\mathrm{CDW}}=(f_{\mathrm{min}}+f_{\mathrm{max}})/2=0.09\,\mathrm{\AA}^{-1}$. Since the CDW peak width is much larger than the REXS momentum resolution, the integration volume is essentially resolution independent: $V(\bm{Q}_{\mathrm{CDW}})=(f_{\mathrm{CDW}}^{2}+f_{\mathrm{res}}^{2})^{3/2}\approx f_{\mathrm{CDW}}^{3}$. Considering data from Fig. 4(B), we find that atomic displacements at 4.95 keV incident energy are limited by $u\leq 0.59\,\mathrm{\AA}$ along the $(100)$ direction and $u\leq 0.1\,\mathrm{\AA}$ along the $(011)$ direction.
While our data set relatively strong lower bounds on absolute intensity limits of charge order with short correlation lengths, the bounds on maximum displacement amplitudes are still relatively weak. The reason here is that short correlation lengths naturally extend over larger regions in momentum space and result, in turn, in relatively small diffraction peak amplitudes ($V(\bm{Q}^{\mathrm{CDW}}_{1,a})\ll V(\bm{G}_{0})$ in Eq. 3). Therefore, we conclude that finite displacements with low correlation lengths are not ruled out by our data.
VII Discussion
@@
Our measurements using resonant X-rays show that the charge modulations earlier observed in STM above the superconducting transition at the surface of UTe2 are absent from the bulk within our detection limits posed by resonant X-ray diffraction. Our measurements were performed at $T=2.2$ K, about four times lower than the highest temperature the CDW in UTe2 is reported to exist [25]. Our results did not reveal diffraction intensity in the vicinity of the CDW wave vectors at either the U M4 or Te L1 absorption edges above the background level.
@@
The coverage of our fine mesh-grid search near the $\bm{q}_{1}$ CDW wave vector rules out the presence of sharp resolution-limited charge-order peaks. Notably, the maximum diffraction intensity of such a CDW peak at the U M edge is at least five orders of magnitude smaller relative to the intensity of the strongest Bragg peaks observed. Atomic displacements of such modulations are four orders of magnitude smaller than the respective atomic units along $(100)$ and $(011)$.
@@
We further investigated reciprocal space with respect to CDW peaks having short correlation lengths in the bulk. Comparing REXS intensities, we find that maximum intensities of such charge order peaks are at least eight orders of magnitude weaker than nearby structural Bragg peaks. We found further that finite atomic displacements of such modulations are not ruled out by our data.
@@
We highlight that neither thermodynamic measurements [29, 30, 31] nor inelastic neutron scattering identified any signatures suggestive of bulk charge order above the superconducting transition [40, 41, 42, 43]. In combination with our results, the emergence of a surface charge-density wave in the normal state provides the most consistent picture. As for the superconducting state, our results point to two possible conclusions: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or these charge modulations are restricted to the surface. Bulk scattering measurements in the superconducting state are required to address this outstanding question.
We note that two additional experimental investigations of UTe2 at low temperature–a non-resonant X-ray diffraction measurement [44] as well as pulse echo and resonant ultrasound measurements [45]–have been reported concurrently with this work. Notably, both studies, performed independently and on samples grown by different groups, are in agreement with our results. Our study is complementary and using resonant X-rays we provide bounds on the structure factor posed detection limits in our experiments.
VIII Materials and Methods
Sample Preparation
@@ VIII Materials and Methods
Resonant Elastic X-ray Scattering
Experiments were carried out in the second experimental hutch EH2 of beamline P09 at the synchrotron source PETRA III [46]. The X-ray diffraction was carried out in a horizontal scattering geometry. The UTe2 sample was cooled using a variable temperature insert in a cryomagnet able to provide magnetic fields up to 14 T. X-rays with incident linear polarization were chosen using phase plates [47].
We use the conventional orthorhombic basis with lattice parameters $a=4.16\,\mathrm{\AA}$, $b=6.13\,\mathrm{\AA}$, and $c=13.97\,\mathrm{\AA}$ for the description of UTe2 crystals. Momentum transfers are denoted by capital bold italic letters, $\bm{Q}$, and given in the reference-frame of the three-dimensional Brillouin zone of UTe2 by means of $\bm{Q}=H\bm{b}_{1}+K\bm{b}_{2}+L\bf{b}_{3}$, where $\bm{b}_{i}=\epsilon_{ijk}2\pi/V\cdot\bm{b}_{j}\times\bm{b}_{k}$. $H$, $K$, $L$ are the Miller indices and $\epsilon_{ijk}$ is the Levi-Civita symbol and $V$ the volume of the conventional unit cell in real-space. Propagation vectors of bulk long-range order are denoted by lowercase bold italic letters, $\bm{q}$. Momentum transfers of structural Bragg peaks are labelled with the letter $\bm{G}$
Characteristic high-symmetry points in the three-dimensional Brillouin zone of UTe2 are denoted by capital letters. Points and coordinates in the two-dimensional reciprocal lattice, used to describe the Fourier transformation of STM images, are given in lowercase letters.
@@
where $A$ denotes the amplitude (or the maximum) and $E$ is a Gaussian profile with full width at half maxima $f_{\bm{d}_{1}}$, $f_{\bm{d}_{2}}$, and $f_{\bm{d}_{3}}$ along the three directions $\bm{d}_{1}=(1,0,0)$, $\bm{d}_{2}=(0,1,1)$, and $\bm{d}_{3}=\bm{d}_{1}\times\bm{d}_{2}$, respectively. Integrated intensity of a Bragg peak is therefore given by $I=\frac{1}{8}\sqrt{\frac{\pi}{\ln(2)}}^{3}\cdot Af_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$. The profile of structural Bragg peaks is essentially restricted by resolution. We therefore define $f_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$ as the experimental resolution volume in momentum space.
Structure factors of UTe2 for X-ray scattering were calculated with scattering amplitudes provided by the International Tables for Crystallography, Ref [32]. Details on these calculations are provided in the Supplementary Information, Text S5.
Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on request.